OpenStudy (anonymous):

Solve (x + 3)(x – 3) = 40. Identify any double roots. A. {3, –3} B. {7, –7} C. {0, 3, –3} D. {0, 7 (double root)}

OpenStudy (anonymous):

Please anyone help!:(

OpenStudy (anonymous):

so FOIL

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

First expand the factored expression and then bring the 40 over to the other side so the whole expression equals 0 and then re-factor to get the x-intercepts, like this:\[(x+3)(x-3)=40 \rightarrow x^2-9=40 \rightarrow x^2-49 = 0 \rightarrow (x-7)(x+7)=0\]\[\implies x = \pm7\]

OpenStudy (anonymous):

@RH

OpenStudy (anonymous):

Therefore, the answer is B. {7, -7} @RH

OpenStudy (anonymous):

yeah {7,-7} i had that also :)

OpenStudy (anonymous):

Thank you both a lot!!!!!!!! :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

Lol medal/fan? @RH =[

OpenStudy (anonymous):

^hahah yeaH?

OpenStudy (anonymous):

Well, you both deserve it, but I am going to be fair and give it to @some_someone because he/she answered first. I am sorry @genius12 . But I thank you both because I understood it quite well.

OpenStudy (anonymous):

thanks :}